# How Transformers Implement Induction Heads: Approximation and Optimization Analysis

### v2

@@ Appendix A Proofs in Section 4
$\begin{pmatrix}x_{s}\\ y_{s}\end{pmatrix}=\begin{pmatrix}x_{s}\\ {\rm softmax}\left(-p^{(1,1)}(s-1-\tau)\right)x_{\tau}\end{pmatrix}.$
@@
$\sum_{s=2}^{L-1}{\rm softmax}(x_{L}^{\top}W^{\star}y_{s})x_{s}$
By Lemma D.1 , for any $L\in\mathbb{N}^{+}$
@@
$=$ $\|\sum_{s=2}^{L-1}{\rm softmax}(x_{L}^{\top}W^{\star}y_{s})x_{s}-\sum_{s=2}^{L-1}{\rm softmax}(x_{L}^{\top}W^{\star}x_{s-1})x_{s}\|_{\infty}$
$\leq$ $\sum_{s=2}^{L-1}|{\rm softmax}(x_{L}^{\top}W^{\star}y_{s})-{\rm softmax}(x_{L}^{\top}W^{\star}x_{s-1})|$
$\leq$ $2\sup_{s}|x_{L}^{\top}W^{\star}y_{s}-x_{L}^{\top}W^{\star}x_{s-1}|$
@@
$\leq$ $2\sum_{i,j}|W_{i,j}^{\star}|\sup_{s}\left\|\left(\sum_{\tau=1}^{s-1}{\rm softmax}\left(-p^{(1,1)}(s-1-\tau)\right)x_{\tau}\right)-x_{s-1}\right\|_{\infty}$
$\leq$ $2\|W^{\star}\|_{1,1}\sup_{s}\left\|\bigg{(}{\rm softmax}\left(-p^{(1,1)}((s-1-\tau))\right)\bigg{)}_{\tau=1}^{s-1}-\bm{e}_{s-1}\right\|_{1}$
$=$ $4\|W^{\star}\|_{1,1}\sup_{s}\left|\frac{1}{\sum_{\tau=0}^{s-2}e^{-p^{(1,1)}\tau}}-1\right|$
$<$ $4\|W^{\star}\|_{1,1}\frac{e^{-p^{(1,1)}}}{1-e^{-p^{(1,1)}}}\leq\mathcal{O}\left(e^{-p^{(1,1)}}\right).$
A.2 Proof of Theorem 4.3
$\textnormal{{IH}}_{n}(X_{L})=\sum_{s=n}^{L-1}{\rm softmax}\Big{(}({x}_{L},\cdots,{x}_{L-n+2})^{\top}W^{\star}({x}_{s-1},\cdots,{x}_{s-n+1})\Big{)}x_{s},$ (14)
Theorem A.2 (Restatement of Theorem 4.3).
Let $\textnormal{{IH}}_{n}$ satisfy Eq. (14). Then for any $q\in\mathbb{N}^{+}$ , there exists a constant $C_{q,n}>0$ and a two-layer $H$-head transformer ${\textnormal{{TF}}}(\cdot)$ (without FFNs), with $D=nd$, such that:
$\sup_{L\in\mathbb{N}^{+}}\left|\!\left|\!\left|\textnormal{{IH}}_{n}-{\textnormal{{TF}}}\right|\!\right|\!\right|_{L,\infty}\leq\frac{C_{q,n}}{H^{q}}.$
@@
$\sum_{s=n}^{L-1}{\rm softmax}\Big{(}({x}_{L},\cdots,\hat{x}_{L-n+2})^{\top}$ $W^{\star}({x}_{s-1},\cdots,\hat{x}_{s-n+1})\Big{)}x_{s}$
@@
$(x_{L}^{\top},$ $0_{D-d}^{\top})^{\top}$
Step I. The first layer. We use 1-st Attn with residual to copy the previous tokens $(x_{s-n+1},\cdots,x_{s-1})$ of each token $x_{s}$. We use $H=\sum_{i=1}^{n-1}H_{i}$ attention heads to realize this step, and the following projection matrices are needed:
$P_{i}:=(0_{d\times id},I_{d\times d},0_{d\times(D-(i+1)d)})\in\mathbb{R}^{d\times D},\quad i=1,\ldots,n-1.$
By lemma D.2, for any rate $q\in\mathbb{N}^{+}$, there exist a constant $C(q)$ and a function
@@
$\|\mathbb{I}\left\{\cdot=i\right\}-\phi_{i}^{\exp}(\cdot)\|_{\ell_{1}(\mathbb{N})}=\sum_{s=i}^{+\infty}|\mathbb{I}\left\{s=1\right\}-\phi^{\exp}(s)|\leq C(q)\frac{e^{0.01(n+1)}}{H_{i}^{n}}$
For $h=\sum_{j=1}^{i-1}H_{j},\ldots,\sum_{j=1}^{i}H_{j}$, we choose parameters as follows
@@
$\left(P_{i}\sum_{h=\sum_{j=1}^{i-1}H_{j}}^{\sum_{j=1}^{i}H_{j}}{\textnormal{{SA}}}^{(1,h)}(X_{L}^{(0)})\right)_{-1}=\sum_{h=\sum_{j=1}^{i-1}H_{j}}^{\sum_{j=1}^{i}H_{j}}\alpha_{h,i}\sum_{s=1}^{L-1}e^{-\beta_{h,i}(s-1)}x_{L-s}.$
We denote $x_{L}^{(1)}:={\textnormal{{SA}}}^{(1)}(X_{L}^{(0)})_{-1}$, then the approximation error of this step is
$\varepsilon_{{\textnormal{{SA}}}}^{(1)}:=\sup_{s}\left\|x_{s}^{(1)}-\begin{pmatrix}x_{s}\\ \vdots\\ x_{s-n+1}\end{pmatrix}\right\|_{\infty}$
$\leq$ $\sup_{s}\sum_{i=1}^{n-1}\left\|P_{i}x_{s}^{(1)}-x_{s-i}\right\|_{\infty}\leq\sup_{s}\sum_{i=1}^{n-1}\|\mathbb{I}\left\{\cdot=i\right\}-\phi_{i}^{\exp}(\cdot)\|_{\ell_{1}(\mathbb{N})}\leq\frac{C_{q,n}}{H^{n}}.$
We denote the output of this step as
$x_{L}^{(1)}:=\begin{pmatrix}x_{L}\\ \hat{x}_{L-1}\\ \vdots\\ \hat{x}_{L-n+1}\end{pmatrix}.$
We choose $H$ large enough so that $\hat{x}_{s-1}^{\top}\in[-2,2]^{d}$.
@@
Then the output of this layer is
$x_{L}^{(2)}=\sum_{s=n}^{L-1}{\rm softmax}\Big{(}({x}_{L},\cdots,\hat{x}_{L-n+2})^{\top}W^{\star}({x}_{s-1},\cdots,\hat{x}_{s-n+1})\Big{)}x_{s}$
According to Lemma D.1,
$\left\|x_{L}^{(2)}-\sum_{s=n}^{L-1}{\rm softmax}\Big{(}({x}_{L},\cdots,{x}_{L-n+2})^{\top}W^{\star}({x}_{s-1},\cdots,{x}_{s-n+1})\Big{)}x_{s}\right\|_{\infty}$
$\leq$ $\sum_{s=n}^{L-1}|{\rm softmax}\Big{(}({x}_{L},\cdots,\hat{x}_{L-n+2})^{\top}W^{\star}({x}_{s-1},\cdots,\hat{x}_{s-n+1})\Big{)}$
$\qquad-{\rm softmax}\Big{(}({x}_{L},\cdots,{x}_{L-n+2})^{\top}W^{\star}({x}_{s-1},\cdots,{x}_{s-n+1})\Big{)}|$
$\leq$ $2\max_{s}|\Big{(}({x}_{L},\cdots,\hat{x}_{L-n+2})^{\top}W^{\star}({x}_{s-1},\cdots,\hat{x}_{s-n+1})\Big{)}$
$\qquad\qquad-\Big{(}({x}_{L},\cdots,{x}_{L-n+2})^{\top}W^{\star}({x}_{s-1},\cdots,{x}_{s-n+1})\Big{)}|$
$\leq$ $2\|W^{\star}\|_{(1,1)}\cdot\varepsilon_{{\textnormal{{SA}}}}^{(1)}.$
Since the above inequality holds for any $L$ and $X_{L}$, we have:
$\sup_{L\in\mathbb{N}^{+}}\left|\!\left|\!\left|\textnormal{{IH}}_{n}-{\textnormal{{TF}}}\right|\!\right|\!\right|_{L,\infty}\leq\frac{\tilde{C}_{q,n}}{H^{q}}.$
A.3 Proof of Theorem 4.4
A.3.1 Approximation Results for FFNs
@@
The well-known universal approximation result for two-layer FNNs asserts that two-layer FNNs can approximate any continuous function [Barron, 1992, 1993, 1994]. Nonetheless, this result lacks a characterization of the approximation efficiency, i.e., how many neurons are needed to achieve a certain approximation accuracy? Extensive pre-existing studies aimed to address this gap by establishing approximation rates for two-layer FFNs. A representative result is the Barron theory [E et al., 2019, 2021, Ma et al., 2020]: any function $f$ in Barron space $\mathcal{B}$ can be approximated by a two-layer FFN with $M$ hidden neurons can approximate $f$ efficiently, at a rate of $\mathcal{O}(\|f\|_{\mathcal{B}}/\sqrt{M})$. This rate is remarkably independent of the input dimension, thus avoiding the Curse of Dimensionality. Specifically, Barron space is defined in as follows:
Definition A.3 (Barron space [E et al., 2019, 2021, Ma et al., 2020]).
@@
$\left\|f-\textnormal{{FFN}}\right\|_{L^{\infty}([0,1]^{d})}\leq\tilde{\mathcal{O}}\left(\frac{\left\|f\right\|_{\mathcal{B}}}{\sqrt{M}}\right).$
A.3.2 Proper Orthogonal Decomposition
Proper orthogonal decomposition (POD) can be viewed as an extension of the matrix singular value decomposition (SVD) applied to functions of two variables. Specifically, for a square integrable function $g:\mathcal{I}\times\mathcal{I}\rightarrow\mathbb{R}$, it has the following decomposition (Theorem 3.4 in  Yarvin and Rokhlin [1998], Theorem VI.17 in Reed and Simon [1980]):

### v3

@@ Appendix C Proofs in Section 3
$\begin{pmatrix}x_{s}\\ y_{s}\end{pmatrix}=\begin{pmatrix}x_{s}\\ (x_{\tau})_{\tau=1}^{s-1}\ {{\rm softmax}}\left(\big{(}-p^{(1,1)}(s-1-\tau)\big{)}_{\tau=1}^{s-1}\right)^{\top}\end{pmatrix}.$
@@
$(x_{s})_{s=2}^{L-1}\ {{\rm softmax}}\Big{(}\big{(}x_{L}^{\top}W^{\star}y_{s}\big{)}_{s=2}^{L-1}\Big{)}^{\top}.$
By Lemma E.1 , for any $L\in\mathbb{N}^{+}$
@@
$=$ $\left\|{(x_{s})_{s=2}^{L-1}\ {\rm softmax}\left(\big{(}x_{L}^{\top}W^{\star}y_{s}\big{)}_{s=2}^{L-1}\right)^{\top}}-{(x_{s})_{s=2}^{L-1}\ {\rm softmax}\left(\big{(}x_{L}^{\top}W^{\star}x_{s-1}\big{)}_{s=2}^{L-1}\right)^{\top}}\right\|_{\infty}$
$\leq$ $\|(x_{s})_{s=2}^{L-1}\|_{\infty,\infty}\left\|{{\rm softmax}\left(\big{(}x_{L}^{\top}W^{\star}y_{s}\big{)}_{s=2}^{L-1}\right)^{\top}}-{{\rm softmax}\left(\big{(}x_{L}^{\top}W^{\star}x_{s-1}\big{)}_{s=2}^{L-1}\right)^{\top}}\right\|_{1}$
$\leq$ $2\sup_{2\leq s\leq L-1}\left|x_{L}^{\top}W^{\star}y_{s}-x_{L}^{\top}W^{\star}x_{s-1}\right|$
@@
$\leq$ $2\sum_{i,j}|W_{i,j}^{\star}|\sup_{s}\left\|(x_{\tau})_{\tau=1}^{s-1}\ {\rm softmax}\left(\left(-p^{(1,1)}(s-1-\tau)\right)_{\tau=1}^{s-1}\right)^{\top}-x_{s-1}\right\|_{\infty}$
$\leq$ $2\|W^{\star}\|_{1,1}\sup_{s}\left\|{\rm softmax}\left(\left(-p^{(1,1)}(s-1-\tau)\right)_{\tau=1}^{s-1}\right)^{\top}-\bm{e}_{s-1}\right\|_{1}$
$=$ $2\|W^{\star}\|_{1,1}\sup_{2\leq s\leq L-1}\frac{\sum_{\tau=1}^{s-2}\exp\left(-p^{(1,1)}(s-1-\tau)\right)}{\sum_{\tau=1}^{s-1}\exp\left(-p^{(1,1)}(s-1-\tau)\right)}$
$\leq$ $2\|W^{\star}\|_{1,1}\lim_{s\to+\infty}\frac{\sum_{\tau=1}^{s-2}\exp\left(-p^{(1,1)}(s-1-\tau)\right)}{\sum_{\tau=1}^{s-1}\exp\left(-p^{(1,1)}(s-1-\tau)\right)}=2\|W^{\star}\|_{1,1}e^{-p^{(1,1)}}.$
C.2 Proof of Theorem 3.3
$\textnormal{{IH}}_{n}(X_{L})=(x_{s})_{s=n}^{L-1}\ {\rm softmax}\Big{(}\big{(}X_{L-n+2:L}^{\top}W^{\star}X_{s-n+1:s-1}\big{)}_{s=n}^{L-1}\Big{)}^{\top},$ (14)
Theorem C.2 (Restatement of Theorem 3.3).
Let $\textnormal{{IH}}_{n}$ satisfy Eq. (14). Then for any $H\in\mathbb{N}^{+}$ and rate $q\in\mathbb{N}^{+}$, there exists a constant $C_{n,q}>0$ and a two-layer $H$-head transformer ${\textnormal{{TF}}}(\cdot)$ (without FFNs), with $D=nd$, such that:
$\sup_{L\in\mathbb{N}^{+}}\left|\!\left|\!\left|\textnormal{{IH}}_{n}-{\textnormal{{TF}}}\right|\!\right|\!\right|_{L,\infty}\leq\left(\frac{C_{n,q}}{H}\right)^{q},$
where $C_{n,q}=\mathcal{O}(nq^{2})$.
@@
$(x_{s})_{s=n}^{L-1}\ {\rm softmax}\Big{(}\big{(}\hat{X}_{L-n+2:L}^{\top}$ $W^{\star}\hat{X}_{s-n+1:s-1}\big{)}_{s=n}^{L-1}\Big{)}^{\top}$
@@
$(x_{L}^{\top},$ $0^{\top})^{\top}$
Additionally, in this proof, the following projection matrices are used:
$P_{i}:=\left(0_{d\times(i-1)d},I_{d\times d},0_{d\times(D-id)}\right)\in\mathbb{R}^{d\times D},\quad i\in[n].$
Step I. The first layer. We use 1-st Attn with residual to copy the previous tokens $(x_{s-n+1},\cdots,x_{s-1})$ of each token $x_{s}$. We use $H=\sum_{i=1}^{n-1}H_{i}$ attention heads to realize this step.
By lemma E.2, for any rate $q\in\mathbb{N}^{+}$, there exists a function
@@
$\|\mathbb{I}\left\{\cdot=i\right\}-\phi_{i}^{\exp}(\cdot)\|_{\ell_{1}(\mathbb{N})}=\sum_{s=i}^{+\infty}|\mathbb{I}\left\{s=1\right\}-\phi^{\exp}(s)|\leq C\frac{A^{q}(q^{2})^{q}e^{0.01(q+1)i}}{H_{i}^{q}},$
where $A,C>0$ are absolute constants.
For $h=\sum_{j=1}^{i-1}H_{j},1+\sum_{j=1}^{i-1}H_{j},\ldots,\sum_{j=1}^{i}H_{j}$, we choose parameters as follows
@@
$\left(P_{i+1}\sum_{h=\sum_{j=1}^{i-1}H_{j}}^{\sum_{j=1}^{i}H_{j}}{\textnormal{{SA}}}^{(1,h)}(X_{L}^{(0)})\right)_{-1}=\sum_{h=\sum_{j=1}^{i-1}H_{j}}^{\sum_{j=1}^{i}H_{j}}\alpha_{h,i}\sum_{s=1}^{L-1}e^{-\beta_{h,i}(s-1)}x_{L-s}.$
We denote the $s$-th output token of the first layer
$x_{s}^{(1)}:={\textnormal{{SA}}}^{(1)}\left(X_{0:s}^{(0)}\right)_{-1},$
Then the approximation error of this step is
$\varepsilon_{{\textnormal{{SA}}}}^{(1)}:=\sup_{s}\left\|x_{s}^{(1)}-X_{s-n+1:s}\right\|_{\infty}\leq\sup_{s}\left(\|x_{s}-x_{s}\|_{\infty}+\sum_{i=1}^{n-1}\left\|P_{i+1}x_{s}^{(1)}-x_{s-i}\right\|_{\infty}\right)$
$=$ $\sup_{s}\sum_{i=1}^{n-1}\left\|P_{i+1}x_{s}^{(1)}-x_{s-i}\right\|_{\infty}\leq\sup_{s}\sum_{i=1}^{n-1}\|\mathbb{I}\left\{\cdot=i\right\}-\phi_{i}^{\exp}(\cdot)\|_{\ell_{1}(\mathbb{N})}$
$\leq$ $C(Aq^{2})^{q}\sum_{i=1}^{n-1}\frac{e^{0.01(q+1)i}}{H_{i}^{q}}.$
Consequently, one detail is to assign the head number $\{H_{i}\}_{i=1}^{n}$ such that the error’s sum $\sum_{i=1}^{n-1}\frac{e^{0.01(q+1)i}}{H_{i}^{q}}$ is as small as possible. Our way is solving the minimization problem
$\min:$ $\sum_{i=1}^{n-1}\frac{e^{0.01(q+1)i}}{H_{i}^{q}}$
${\rm s.t.}$ $\sum_{i=1}^{n-1}H_{i}=H,$
which suggests that we should choose the head number:
$H_{i}=\frac{e^{0.01i}}{\sum_{j=1}^{n-1}e^{0.01j}},i\in[n-1].$
Thus, we obtain the bound
$\varepsilon_{{\textnormal{{SA}}}}^{(1)}\leq\frac{C(Aq^{2})^{q}}{H^{q}}\left(\sum_{i=1}^{n-1}e^{0.01i}\right)^{q}\leq C\left(\frac{Aq^{2}ne^{0.01n}}{H}\right)^{q}.$
Noticing $n<100$, we have:
$\varepsilon_{{\textnormal{{SA}}}}^{(1)}\leq C\left(\frac{eAq^{2}n}{H}\right)^{q}.$
We focus on the case of large $H$:
$H\geq CAenq^{2},$
which ensures $\varepsilon_{{\textnormal{{SA}}}}^{(1)}\leq 1$. Therefore, $x_{s}^{(1)}\in[-2,2]^{D}$:
$\left\|x_{s}^{(1)}\right\|_{\infty}\leq\sup_{s}\left\|x_{s}^{(1)}-X_{s-n+1:s}\right\|_{\infty}+\left\|X_{s-n+1:s}\right\|_{\infty}\leq\varepsilon_{{\textnormal{{SA}}}}^{(1)}+1\leq 2,\quad\forall s.$
@@
For simplicity, we use the following notations:
$\hat{X}_{s-n+1:s-1}:=W_{Q}=\begin{pmatrix}P_{2}{x}_{s}^{(1)}\\ \vdots\\ P_{n}{x}_{s}^{(1)}\end{pmatrix},\ \text{ for }n\leq s\leq L-1;\quad\quad\hat{X}_{L-n+2:L}:=\begin{pmatrix}P_{1}{x}_{L}^{(1)}\\ \vdots\\ P_{n-1}{x}_{L}^{(1)}.\end{pmatrix}.$
Then the output of the second layer is
$x_{L}^{(2)}=(x_{s})_{s=n}^{L-1}\ {\rm softmax}\Big{(}\big{(}\hat{X}_{L-n+2:L}^{\top}W^{\star}\hat{X}_{s-n+1:s-1}\big{)}_{s=n}^{L-1}\Big{)}^{\top}$
By using these bounds and Lemma E.1,
$\left\|x_{L}^{(2)}-(x_{s})_{s=n}^{L-1}\ {\rm softmax}\Big{(}\big{(}X_{L-n+2:L}^{\top}W^{\star}X_{s-n+1:s-1}\big{)}_{s=n}^{L-1}\Big{)}^{\top}\right\|_{\infty}$
$\leq$ $\sum_{s=n}^{L-1}\Bigg{|}{\rm softmax}\Big{(}\big{(}\hat{X}_{L-n+2:L}^{\top}W^{\star}\hat{X}_{s-n+1:s-1}\big{)}_{s=n}^{L-1}\Big{)}^{\top}-{\rm softmax}\Big{(}\big{(}X_{L-n+2:L}^{\top}W^{\star}X_{s-n+1:s-1}\big{)}_{s=n}^{L-1}\Big{)}^{\top}\Big{)}\Bigg{|}$
$\leq$ $2\max_{s}\left|\hat{X}_{L-n+2:L}^{\top}W^{\star}\hat{X}_{s-n+1:s-1}-{X}_{L-n+2:L}^{\top}W^{\star}{X}_{s-n+1:s-1}\right|$
$\leq$ $2\Big{(}\max_{s}\left|\left(\hat{X}_{L-n+2:L}-{X}_{L-n+2:L}\right)^{\top}W^{\star}\hat{X}_{s-n+1:s-1}\right|+\max_{s}\left|X_{L-n+2:L}^{\top}W^{\star}\left(\hat{X}_{s-n+1:s-1}-{X}_{s-n+1:s-1}\right)\right|\Big{)}$
$\leq$ $2\Big{(}2\|W^{\star}\|_{(1,1)}\left\|\hat{X}_{L-n+2:L}-{X}_{L-n+2:L}\right\|_{\infty}+\|W^{\star}\|_{1,1}\max_{s}\left\|\hat{X}_{s-n+1:s-1}-{X}_{s-n+1:s-1}\right\|_{\infty}\Big{)}$
$\leq$ $6\|W^{\star}\|_{1,1}\cdot\varepsilon_{{\textnormal{{SA}}}}^{(1)}.$
Since the above inequality holds for any $L$ and $X_{L}$, we obtain our bound:
$\sup_{L\in\mathbb{N}^{+}}\left|\!\left|\!\left|\textnormal{{IH}}_{n}-{\textnormal{{TF}}}\right|\!\right|\!\right|_{L,\infty}\leq 6C\|W^{\star}\|_{(1,1)}\left(\frac{6eAnq^{2}}{H}\right)^{q}\leq\left(\frac{6C\|W^{\star}\|_{1,1}eAnq^{2}}{H}\right)^{q},$
For the remained case, $H<CAenq^{2}$, we can simply choose TF with all 00 parameters. Then the approximation error can be trivially bounded by:
$\sup_{L\in\mathbb{N}^{+}}\left|\!\left|\!\left|\textnormal{{IH}}_{n}-{\textnormal{{TF}}}\right|\!\right|\!\right|_{L,\infty}=\sup_{L\in\mathbb{N}^{+}}\left|\!\left|\!\left|\textnormal{{IH}}_{n}-0\right|\!\right|\!\right|_{L,\infty}\leq 1.$
Then, the two cases can be unified by:
$\sup_{L\in\mathbb{N}^{+}}\left|\!\left|\!\left|\textnormal{{IH}}_{n}-{\textnormal{{TF}}}\right|\!\right|\!\right|_{L,\infty}\leq\begin{cases}6C\|W^{\star}\|_{1,1}\left(\frac{eAnq^{2}}{H}\right)^{q},\ H\geq CAenq^{2}\\ 1,\ \text{otherwise}\end{cases}\leq\left(\frac{\left(6\|W^{\star}\|_{1,1}+1\right)CeAnq^{2}}{H}\right)^{q}.$
In the theorem, we can choose $C_{n,q}=\left(6\|W^{\star}\|_{1,1}+1\right)CeAnq^{2}$, which satisfies $C_{n,q}\lesssim nq^{2}$. ∎
C.3 Proof of Theorem 3.4
C.3.1 Approximation Results for FFNs
@@
The well-known universal approximation result for two-layer FNNs asserts that two-layer FNNs can approximate any continuous function (Barron, 1992; 1993; 1994). Nonetheless, this result lacks a characterization of the approximation efficiency, i.e., how many neurons are needed to achieve a certain approximation accuracy? Extensive pre-existing studies aimed to address this gap by establishing approximation rates for two-layer FFNs. A representative result is the Barron theory (E et al., 2019; 2021; Ma et al., 2020): any function $f$ in Barron space $\mathcal{B}$ can be approximated by a two-layer FFN with $M$ hidden neurons can approximate $f$ efficiently, at a rate of $\mathcal{O}(\|f\|_{\mathcal{B}}/\sqrt{M})$. This rate is remarkably independent of the input dimension, thus avoiding the Curse of Dimensionality. Specifically, Barron space is defined in as follows:
Definition C.3 (Barron space (E et al., 2019; 2021; Ma et al., 2020)).
@@
$\left\|f-\textnormal{{FFN}}\right\|_{L^{\infty}[-2,2]}\leq\mathcal{O}\left(\frac{\left\|f\right\|_{\mathcal{B}}\sqrt{\log M}}{\sqrt{M}}\right).$
C.3.2 Proper Orthogonal Decomposition
Proper orthogonal decomposition (POD) can be viewed as an extension of the matrix singular value decomposition (SVD) applied to functions of two variables. Specifically, for a square integrable function $g:\mathcal{I}\times\mathcal{I}\rightarrow\mathbb{R}$, it has the following decomposition (Theorem 3.4 in  Yarvin and Rokhlin (1998), Theorem VI.17 in Reed and Simon (1980)):